# EndoChat: Grounded Multimodal Large Language Model for Endoscopic Surgery

### v1

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have demonstrated their immense potential in computer-aided diagnosis and decision-making. In the context of robotic-assisted surgery, MLLMs can serve as effective tools for surgical training and guidance. However, there is still a lack of MLLMs specialized for surgical scene understanding in clinical applications. In this work, we introduce EndoChat to address various dialogue paradigms and subtasks in surgical scene understanding that surgeons encounter. To train our EndoChat, we construct the Surg-396K dataset through a novel pipeline that systematically extracts surgical information and generates structured annotations based on collected large-scale endoscopic surgery datasets. Furthermore, we introduce a multi-scale visual token interaction mechanism and a visual contrast-based reasoning mechanism to enhance the model’s representation learning and reasoning capabilities. Our model achieves state-of-the-art performance across five dialogue paradigms and eight surgical scene understanding tasks. Additionally, we conduct evaluations with professional surgeons, most of whom provide positive feedback on collaborating with EndoChat. Overall, these results demonstrate that our EndoChat has great potential to significantly advance training and automation in robotic-assisted surgery.
Introduction
@@
Robot-assisted surgery (RAS) offers unprecedented opportunities to enhance surgical precision, minimize patient trauma, and shorten postoperative recovery times [1]. However, the effective application of this technology places significant demands on the skills of surgeons, particularly in mastering the operation of robotic systems during procedures. To ensure surgical safety and efficacy, surgeons must undergo rigorous training to acquire and refine the core skills required for robotic operation [2, 3]. To improve the efficiency of this training process, various simulator-based surgical platforms [4] have been developed. However, when trainees encounter challenges during training, they often require immediate feedback and guidance from professional surgeons to resolve doubts or correct mistakes. Unfortunately, professional surgeons typically face significant time constraints due to their heavy workload in clinical, teaching, and research responsibilities, making it difficult for them to provide consistent, real-time support during training [5, 6]. As a result, there is a pressing need for technological solutions that can deliver flexible, real-time, and efficient support in surgical training.
@@
Recently, artificial intelligence (AI)-based dialogue systems with structured Visual Question Answering (VQA) have been introduced into surgical training [7, 8, 9, 10, 11, 12, 13, 14, 6, 15, 16, 17]. These systems analyze visual data from surgical scenarios to address trainees’ questions. As the first work to propose a question-answering (QA) model specifically designed for surgical scenarios, Surgical-VQA represents key elements in surgical scenes—such as instruments, tissues, tools, and spatial positions—through textual descriptions [6]. Using the VisualBERT [18] framework, it integrates multimodal representations of text and images and generates corresponding answers through a decoder. Subsequent works, such as Surgical-VQLA, enhance this approach by incorporating bounding box outputs at the decoder level, providing explicit visual localization to assist surgeons [7, 9, 12, 17]. Later developments explored different network designs or inference paradigms. However, these models often rely on structured VQA datasets, which are typically designed and trained for specific tasks. This reliance limits their ability to dynamically adapt to the wide range of questions posed by trainees [19, 20]. Moreover, most of these approaches are based on Encoder-Decoder architectures, which require explicitly defined input/output formats. This rigidity makes them less flexible for handling highly open-ended generative tasks and limits their scalability. When applied beyond their trained scope, these models often show significant performance drops, making them poorly suited for the complexity of diverse surgical scenarios [21, 22]. When trainees ask open-ended questions, existing VQA systems lack the flexibility and contextual understanding needed for such interactions. As a result, they struggle to handle open-ended questions or complex multi-turn dialogues, significantly limiting their usefulness in surgical training.
@@
Recently, medical Multimodal Large Language Models (MLLMs) are emerging as a promising solution, offering significant potential through large-scale pretraining to perform complex reasoning and understanding across tasks [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. Specifically, MLLMs can extract information from multimodal data in surgical scenarios and perform advanced reasoning. Unlike structured question-answering systems, MLLMs are capable of processing unstructured and complex contextual information. For example, trainees can ask open-ended questions in natural language, and MLLMs can generate targeted responses by leveraging their pre-trained knowledge and multimodal reasoning capabilities. These powerful natural language processing abilities enable MLLMs to handle multi-turn conversations and dynamically adjust responses based on context. This interaction is akin to receiving guidance from professional surgeons, significantly enhancing the training experience. Overall, MLLMs have the potential to partially replace real-time guidance from professional surgeons by simulating their knowledge and decision-making abilities. This could address the limitations of current training solutions and alleviate the burden on surgeons who are constrained by their demanding clinical schedules, ultimately improving the efficiency and quality of surgical training [26, 40, 41].
Some prior works have explored the capabilities of MLLMs in surgical scenarios [42, 43, 44, 45, 46]. For instance, LLaVA-Surg collects and annotates open-source videos and datasets, then enhances the model’s conversational abilities through instruction tuning [42]. GP-VLS, on the other hand, unifies various tasks in surgical scenarios into a question-answering framework, modeling and expressing the results of different downstream surgical tasks in textual form [45]. However, we aim to further develop MLLMs to handle diverse queries from trainees in real-world scenarios. For example, a trainee might ask a brief question about a specific target, which does not require any redundant information. Therefore, Single Phrase QA is designed to provide concise, accurate, and direct responses relevant to surgical contexts. When a trainee requests a detailed explanation of an entire surgical image, Detailed Description, on the other hand, offers surgeons comprehensive explanations of all sub-tasks present within the current surgical scene. For routine queries about image content, Visual QA delivers context-aware answers by combining user queries with image information. When a trainee needs a focused conversation about a specific region, Region Based QA focuses on targeted responses for specific areas of interest, while Grounding QA generates bounding box descriptions based on user-provided prompts. These five distinct conversational paradigms encompass the majority of scenarios encountered in natural language dialogues [47]. By constructing them, we aim to design a surgical MLLM that can respond appropriately in different cases, creating a system that is better suited for human interaction. This would provide more practical and effective support for surgical training and education.

### v2

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have demonstrated their immense potential in computer-aided diagnosis and decision-making. In the context of robotic-assisted surgery, MLLMs can serve as effective tools for surgical training and guidance. However, there is still a deficiency of MLLMs specialized for surgical scene understanding in endoscopic procedures. To this end, we present EndoChat MLLM to address various dialogue paradigms and subtasks in understanding endoscopic procedures. To train our EndoChat, we construct the Surg-396K dataset through a novel pipeline that systematically extracts surgical information and generates structured annotations based on large-scale endoscopic surgery datasets. Furthermore, we introduce a multi-scale visual token interaction mechanism and a visual contrast-based reasoning mechanism to enhance the model’s representation learning and reasoning capabilities. Our model achieves state-of-the-art performance across five dialogue paradigms and seven surgical scene understanding tasks. Additionally, we conduct evaluations with professional surgeons, who provide positive feedback on the majority of conversation cases generated by EndoChat. Overall, these results demonstrate that EndoChat has the potential to advance training and automation in robotic-assisted surgery. Our dataset and model are publicly available at https://github.com/gkw0010/EndoChat.
keywords:
Multimodal large language model, Endoscopic surgery, Surgical scene understanding, Dialogue paradigm
[…]
1 Introduction
@@
Robot-assisted surgery (RAS) offers unprecedented opportunities to enhance surgical precision, minimize patient trauma, and shorten postoperative recovery times [54, 55, 34]. However, the effective application of this technology places significant demands on the skills of surgeons, particularly in mastering the operation of robotic systems during procedures [2, 74]. To ensure surgical safety and efficacy, surgeons must undergo rigorous training to acquire the core skills required for robotic operation [20, 6]. To improve the efficiency of this training process, various simulator-based surgical platforms [53] have been developed. However, when trainees encounter challenges during training, they often require immediate feedback and guidance from professional surgeons to resolve doubts or correct mistakes. Unfortunately, professional surgeons typically face significant time constraints due to their heavy workload in clinical, teaching, and research responsibilities, making it difficult for them to provide consistent, real-time support during training [64, 63]. As a result, there is a pressing need for technological solutions that can deliver flexible, real-time, and efficient support in surgical training.
@@
Recently, artificial intelligence (AI)-based dialogue systems with structured Visual Question Answering (VQA) have been introduced into surgical training [63, 9, 28]. These systems analyze visual data from surgical scenarios to address trainees’ questions. However, their reliance on simple structured and object presence-based VQA datasets, typically trained for specific tasks, limits their ability to dynamically adapt to the wide range of questions posed by trainees [57, 47]. Moreover, most of these approaches are based on encoder-decoder architectures, which require explicitly defined input/output formats. This rigidity makes them less flexible for handling highly open-ended generative tasks and limits their scalability. When applied beyond their designed scope, these models often show significant performance drops, making them poorly suited for the complexity of diverse surgical scenarios [22, 26]. When trainees ask open-ended questions, existing VQA systems lack the flexibility and contextual understanding needed for such interactions. As a result, they struggle to handle open-ended questions or complex multi-turn dialogues, significantly limiting their usefulness in surgical training.
@@
Currently, medical Multimodal Large Language Models (MLLMs) are emerging as a promising solution, offering significant potential through large-scale pretraining to perform complex reasoning and understanding across tasks [40, 15, 19, 48]. Specifically, MLLMs can extract information from multimodal data in surgical scenarios and perform advanced reasoning. Unlike structured question-answering systems, MLLMs are capable of processing unstructured and complex contextual information. For example, trainees can ask open-ended questions in natural language, and MLLMs can generate targeted responses by leveraging their pre-trained knowledge and multimodal reasoning capabilities. These powerful natural language processing abilities enable MLLMs to handle multi-turn conversations and dynamically adjust responses based on context. This interaction is akin to receiving guidance from professional surgeons, significantly enhancing the training experience. Overall, MLLMs have the potential to partially replace guidance from professional surgeons by simulating their knowledge and decision-making abilities. This can address the limitations of current training solutions and alleviate the burden on surgeons who are constrained by demanding clinical schedules, ultimately improving the efficiency and quality of surgical training [15, 23, 29].
Surgical MLLMs have attracted significant attention from researchers [42, 35, 75]. However, existing surgical MLLMs primarily focus on one of three approaches: leveraging the powerful capabilities of MLLMs to enhance question-answering performance [30], extending models for specific predefined tasks through instruction tuning [75], or using web-sourced data to construct general-purpose descriptive and conversational datasets [42]. These methods face two major challenges: (i) In real-world scenarios, the queries from trainees are highly diverse. Solely relying on predefined query formats or generic captioning is insufficient to handle the variety of queries from trainees in practical settings. (ii) Many current MLLM applications rely on pre-trained visual encoders to extract visual features [42, 50]. However, due to the domain gap, semantic information from general-purpose visual scenes cannot be directly applied to surgical scenarios. This leads to an inadequate understanding of visual information and the occurrence of reasoning hallucinations.
To this end, we address the above challenges from two perspectives: constructing open-ended, knowledge-required surgical vision-language datasets and developing vision-enhanced MLLM models tailored for surgical scenarios. Firstly, we develop Surg-396K, a surgical multimodal instruction dataset specifically for surgical contexts. We systematically extract surgical attribute information from three public datasets and generate instruction-tuning data using diverse conversational templates. To better simulate real-world scenarios, as presented in Table 1, we propose five distinct conversational paradigms to capture the majority of natural language dialogue scenarios and define seven attribute-specific, surgery-related downstream tasks to ensure comprehensive coverage of surgical scene understanding. As a result, the MLLM trained on Surg-396K is better equipped for surgeon-system interaction, enabling it to respond effectively in surgical contexts and provide practical, reliable support for surgical training and education. Secondly, in the model architecture of MLLMs, we design the Mixed Visual Token Engine (MVTE) to extract visual information at multiple scales for better vision-language alignment in surgical contexts. Unlike traditional frameworks that directly use pre-trained Vision Transformers (ViTs) to extract visual tokens, MVTE uses multiple visual towers to extract, interact, and fuse visual tokens, improving the visual information extraction before aligning with text. Furthermore, to minimize model hallucinations, we propose a visual contrast-based approach that compares outputs generated from both original and distorted visual inputs. This method refines the token selection process by applying adaptive plausibility constraints, ensuring better consistency between visual inputs and language outputs in complex endoscopic surgery scenarios